OpenStudy (anonymous):

There is a Serie \(\sum_{n=1}^{\infty} \frac{x^{2n}}{n^{2}}\) Convergence radius \(r = 1\) is convergent in both endpoints Set \(y=x^{2}\) Look for convergence radius \(r\) of Serie \(\sum_{n=1}^{\infty} \frac{y^{n}} {n^{2}}\). Let \(a_{n} = \frac{1}{n^{2}}\) ... I cant understand why or how we can take \(a_{n} = \frac{1}{n^{2}}\)

OpenStudy (anonymous):

also my problem is with \[a_{n} = \frac{1}{n^{2}}\] why we can take it instead of\[\frac{y^{n}}{n^{2}}\]

OpenStudy (zzr0ck3r):

a_n = (1/n)^2

OpenStudy (blockcolder):

If you examine the definition of a power series, it says something about a series of the ff. form: \[\sum c_nx^n\] So in this case, you ignore the variable raised to n (which is y^n) and consider the remaining expression in n.

OpenStudy (zzr0ck3r):

I'm a little confused by the question

OpenStudy (anonymous):

hmm i am confusing in instead of \[y^{n}\] why we write 1

OpenStudy (anonymous):

@blockcolder can you tell me please in other words how we can get \[1\] instead of \[y^{n}\]

OpenStudy (blockcolder):

You don't replace \(y^n\) by 1. \[\frac{y^n}{n^2}=y^n\cdot \frac{1}{n^2}\] As I said previously, you ignore the part with the variable raised to n, and just let the rest be \(a_n\). Another example to make it clear: \[\sum \frac{x^n}{n3^n}=\sum x^n \cdot \frac{1}{n3^n}\] so you let \(a_n=\frac{1}{n3^n}\).

OpenStudy (anonymous):

ok thank you very much i think this gonna helps me

OpenStudy (anonymous):

sure i get it @blockcolder, believe me you will be very good professour insAllah ;)

OpenStudy (blockcolder):

Thanks for the complement. :D